Ammonia is an important compound and useful to industries such as agricultural, pharmaceuticals, textile, plastics, petroleum, mining, and commercial. In the ammonia production, Haber process is utilized and it is shown as, N2(g) + 3H2(g) ➡️ 2NH3(g).



If there are 10.0 moles of ammonia produced, how many moles of nitrogen gas and hydrogen gas react?


1
Expert's answer
2021-11-01T06:32:28-0400

According to the reaction:

Two moles of NH3 are produced from one mole of N2

"n(N_2) = \frac{1}{2}n(NH_3) = \frac{1}{2} \times 10.0 = 5.0 \;moles"

Two moles of NH3 are produced from three moles of H2

"n(H_2) = \frac{3}{2}n(NH_3) = \frac{3}{2} \times 10.0 = 15.0 \;moles"